# The Energy Spectrum of Kaon from Lattice QCD: A Statistical and Theoretical Comparison

### v1

@@ Abstract
This study presents the analysis of data related to the two-point function of kaon generated from lattice QCD simulations. Using gauge configurations of twisted-mass fermions, we obtain the correlation functions for six values of momentum for the kaon between 0 and 2GeV (both inclusive), we use statistical techniques such as jackknife resampling to derive the energy of the particle. The lattice results are compared to the continuum dispersion relation for the particle, to assess systematic uncertainties in the lattice data. We establish consistency with theory by comparing our results with the theoretical predictions.
keywords:
[…]
[d]organization=Indian Institute of Technology Madras, city=Madras, country=India
[…]
1 Introduction
The strong force is one of the four fundamental forces in nature and is described by the theory of Quantum
Chromodynamics (QCD). Gluons are the force carriers, and quarks can interact via this force. Kaons are
among the many hadrons that arise from the strong force. They are further classified as mesons (particles
made of a quark and anti-quark pair). They contain one strange quark (or antiquark) and an up or down
quark (or antiquark), which leads to four types of kaons (Particle Data Group (2020)).
We cannot use our perturbative tools to study the structure of hadrons due to the large strong coupling of the
force in the low energy region (confinement) (Simonov (2019) Shifman (2010)Gross et al. (2023)). Another complication is the fact there are infinite degrees of
freedom, that is quarks and gluons (van Baal (1998)). These degrees of freedom manifest themselves physically in the QCD
vacuum that consists of sea quarks and gluons. We define a finite volume hypercubic grid of space-time
points (lattice) (Gross et al. (2023)). Since the lattice is defined in a finite volume, there are finite degrees of freedom, but,
nevertheless, are of the order of billions.
@@ 1 Introduction
One of the key points that differs QCD from Quantum Electrodynamics (QED) is that gauge invariance implies that the Lagrangian contains terms involving three or four gluon fields: in contrast to the photons, which interact among themselves only via the exchange of charged particles, the gluons would interact even if quarks did not exist (Gross et al. (2023)).
To convert to the lattice QGT formulation, we introduce the spacing ’’ between the points of a hypercubic grid, so the lattice field is defined on a discrete set of points in and .
On a discrete lattice, the derivatives can be replaced by finite difference methods. Combining a rotation to Euclidean space and discretization, we can write the path integral as:
where is the field at each site, and is,
The field is defined on a 4-D Euclidean domain, and we can observe that the partition function looks like that of a statistical mechanical system. This allows us to use Monte Carlo sampling techniques. The observables are calculated as expectation values ( Montvay and Münster (1994) Gross et al. (2023)). Their values change with lattice spacing, but they are usually related to our physical units by simple mathematical relations.
@@ 2 METHODOLOGY
We shall now pursue Lattice QCD more descriptively. It should be noted that the lattice description of Quantum field theories allows us to regularize them without the use of a perturbative approach, the lattice spacing acts as a cutoff for spacetime separation and equivalently for momentum (Hari Dass (1985)). However, this lattice has no direct physical correspondence of its own at a finite ; we are eventually interested in obtaining observables at the continuum limit of the theory i.e., (Wilson (1974) Montvay and Münster (1994) Gross et al. (2023)). However, we have not encountered any explicit dependence on lattice spacing- in the theory; this problem is readily solved by looking at an analogous lattice description of a classical statistical system. This tells us that such a lattice formulation requires certain physical parameters analogous to Temperature() in the classical regime, which allows us to determine all the measurable in terms of other physical parameters.
2.1 How - What - Why ?
The structure of lattice action projects the physical picture of the system onto the lattice, such simulation scheme demands that quarks (or antiquarks) are placed at the lattice points which are linked through the gluonic field (Wilson (1974) Montvay and Münster (1994)). The usual calculation of mass () for such a system requires the creation and annihilation of particle (kaon) inside the hypercubic lattice at time and , the best way to extract the energy of the particle for a certain momentum state (Durr et al. (2008)) is through using Two-point correlation function (Durr et al. (2008) Montvay and Münster (1994)).
@@ 2.1 How - What - Why ?
2.2 Jackknife Resampling
Jackknife resampling (Wu (1986)) is extensively used by the lattice practitioners, it is a ready-to-go tool used for error estimation and information gathering, using a small dataset. Which is what we deal with in lattice QCD systems. An alternative approach is the bootstrap analysis (Efron (1992)), which we do not discuss here. Regarding jackknife, the basic steps are:
We analyzed six values of the particle momentum, that is 0, 0.41, 0.83, 1.24, 1.66, 2.07 GeV. The statistics for each of these momenta is 1 198, 9 584, 9 584, 9 584, 28 752, 239 600, respectively. Also, in each case we have positive and negative values for the momentum. The simulation involves the following steps:
2.3 Mean value of 2pt functions for positive and negative momentum direction and Data Organisation:
We average the real part of two-point functions corresponding to the same magnitude of momentum in the z-direction at each time step. This includes averaging the functions for positive and negative z-direction while keeping the x and y directions constant. Then processed the files by separating the averaged real values of the two-point functions based on time. The data is then organized into a dictionary indexed by time values.
2.4 Data Averaging and time averaging:
Each Dictionary had 41 datasets. We averaged the values associated with the same time value of datasets having the same value of total momentum. Then using time reversal symmetry we averaged the 2pt values at time values (i) and (64-i) for i from 1 to 31. The values for time values 0 and 32 remain unchanged. This step is necessary to obtain a single plateau in the effective energy vs. time graph.
2.5 Jackknife Analysis:
We performed Jackknife resampling on the averaged data to calculate the Jackknife mean and standard deviation for each time step. This helps in estimating the uncertainty of the results.
2.6 Energy Calculation and Identification of Plateau:
We computed the ratio of two-point functions at consecutive time steps and took the natural logarithm to form the energy arrays. After plotting the vs graph we observed a plateau-like region, selected the range in terms of time step[] and mark it as plateau region.
2.7 Calculation of Plateau Energy and Jackknife Error:
Then we calculated the final value of effective energy () at that particular value of momentum using the following formula:
@@ 2.7 Calculation of Plateau Energy and Jackknife Error:
2.8 Comparison with Dispersion Relation:
Next, we compare the obtained results with the continuum dispersion relation using the results of various momentum analyses, applying Einstein’s special theory of relativity. In natural units:
where is the lattice unit.
@@ 3 Results and Conclusion
Following the already mentioned simulation, two representative plateau are shown below [ 1] for the momentum 0.83 GeV and 1.24 GeV (where the x and y component is 0) and momentum (where the resultant of x and y component is ) .
Plateau fitted values for different combinations of momenta boost are given below in the table [ 1]. Now we compare our computationally generated data (through Lattice QCD) with its theoretical counterpart using the dispersion relation [ 2].
From the above plot [ 2] we clearly see that our data clearly agrees with its theoretical counterpart so we can say that Lattice QCD can be one of the promising approach to probe the kaons in lower as well as for higher energy spectrum.
| Momentum | ||
| Configuration | ||
| Effective Energy | Error | |
| 0.2503897571908317 | ||
| 0.31812734594752745 | ||
| 2 | 0.3732724854969631 | |
| 3 | 0.4247528925508162 | |
| 0.463223332509825 | ||
| 5 | 0.5085321794457005 | |
| 6 | 0.5438130645474166 | |
| 8 | 0.5980768744904814 | |
| 0.6356445374137877 | ||
| 10 | 0.6668746069233606 | |
| 11 | 0.69971653807267 | |
| 12 | 0.7251039340525522 | |
| 13 | 0.747680677161704 | |
| 14 | 0.7759013508398236 | |
| 0.8289985937907043 | ||
| 17 | 0.854196899649006 | |
| 18 | 0.8759581328257354 | |
| 19 | 0.8991794712094623 | |
| 20 | 0.91539787219901 | |
| 21 | 0.9344281530391869 | |
| 24 | 0.9972232821194545 | |
| 1.00780980951629 | ||
| 27 | 1.0456348117309648 |
Acknowledgements
We thank Dr. Martha Constantinou for mentoring us throughout the project. We also thank Josh Miller and Isaac Anderson for their advice on technical aspects. M.C. acknowledges financial support from the U.S. Department of Energy, Office of Nuclear Physics, Early Career Award under Grant No. DE-SC0020405, and Grant No. DE-SC0025218. Interactions within the Quark-Gluon Tomography (QGT) Topical Collaboration, funded by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, with Award DE-SC0023646, have benefited this work. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award NP-ERCAP0027642, as well as NERSC award ALCC- ERCAP0030652. Also, some of the data used in this work have been produced using a computational award from Oak Ridge Leadership Computing Facility through the Summit Plus program (project: NPH160).
4 Data and code availability
Data and code will be made available on request.

### v2

# Arunangshu Bora1Mentor: Dr. Martha Constantinou

@@ Abstract
This study presents the analysis of data related to the two-point function of kaon generated from
lattice QCD simulations. Using gauge configurations of twisted-mass fermions, we obtain the correlation
functions for 6 values of momentum for the kaon between 0 and 2GeV(both inclusive), we use statistical
techniques such as jackknife resampling to derive the energy of the particle. The lattice results are
compared to the continuum dispersion relation for the particle, to assess systematic uncertainties in
the lattice data. We establish consistency with theory by comparing our results with the theoretical
predictions.
The strong force is one of the four fundamental forces in nature and is described by the theory of Quantum Chromodynamics (QCD). Gluons are the force carriers, and quarks can interact via this force. Kaons are among the many hadrons that arise from the strong force. They are further classified as mesons (particles made of a quark and anti-quark pair). They contain one strange quark (or antiquark) and an up- or down quark (or antiquark), which leads to four types of kaons. We cannot use our perturbative tools to study the structure of hadrons due to the large strong coupling of the force in the low energy region (confinement). Another complication is the fact there are infinite degrees of freedom, that is quarks and gluons. These degrees of freedom manifest themselves physically in the QCD vacuum that consists of sea quarks and gluons. We define a finite volume hypercubic grid of space-time points (lattice) [[1]]. Since the lattice is defined in a finite volume, there are finite degrees of freedom, but, nevertheless, are of the order of billions.
@@
One of the key points that differs QCD from Quantum Electrodynamics (QED) is that gauge invariance implies that the Lagrangian contains terms involving three or four gluon fields: in contrast to the photons, which interact among themselves only via the exchange of charged particles, the gluons would interact even if quarks did not exist [[1]].
To convert to the lattice QGT formulation, we introduce the spacing ‘’ between the points of a hypercubic grid, so the lattice field is defined on a discrete set of points in and . On a discrete lattice, the derivatives can be replaced by finite difference methods. Combining a rotation to Euclidean space and discretization, we can write the path integral as:
where is the field at each site, and, is,
The field is defined on a 4-D Euclidean domain, and we can observe that the partition function looks like that of a statistical mechanical system. This allows us to use Monte Carlo sampling techniques. The observables are calculated as expectation values. Their values change with lattice spacing, but they are usually related to our physical units by simple mathematical relations.
@@ 2 METHODOLOGY
We shall now pursue Lattice QCD more descriptively. It should be noted that the lattice description of Quantum field theories allows us to regularize them without the use of a perturbative approach, the lattice spacing acts as a cutoff for spacetime separation and equivalently for momentum.[[2]] However, this lattice has no direct physical correspondence of its own at a finite ; we are eventually interested in obtaining observables at the continuum limit of the theory i.e., .
However, we have not encountered any explicit dependence on lattice spacing- in the theory; this problem is readily solved by looking at an analogous lattice description of a classical statistical system. This tells us that such a lattice formulation requires certain physical parameters analogous to Temperature in the classical regime, which allows us to determine all the measurable in terms of other physical parameters.
How - What - Why ? The structure of lattice action projects the physical picture of the system onto the lattice, such simulation scheme demands that quarks (or antiquarks) are placed at the lattice points which are linked through the gluonic field. The usual calculation of mass for such a system requires the creation and annihilation of particle (kaon) inside the hypercubic lattice at time and , the best way to extract the energy of the particle for a certain momentum state is through using Two-point correlation function.
@@
Jackknife resampling is extensively used by the lattice practitioners, it is a ready-to-go tool used for error estimation and information gathering, using a small dataset. Which is what we deal with in lattice QCD systems. An alternative approach is the bootstrap analysis, which we do not discuss here. Regarding jackknife, the basic steps are:
3 PROCEDURE
We analyzed six values of the particle momentum, that is .The statistics for each of these momenta is 1 198, 9 584, 9 584, 9 584, 28 752, 239 600, respectively. Also, in each case we have positive and negative values for the momentum. The procedure involves the following steps:
1.
Mean value of 2pt functions for positive and negative momentum direction and Data Organisation: We average the real part of two-point functions corresponding to the same magnitude of momentum in the z-direction at each time step. This includes averaging the functions for positive and negative z-direction while keeping the x and y directions constant. Then processed the files by separating the averaged real values of the two-point functions based on time. The data is then organized into a dictionary indexed by time values.
2.
Data Averaging and time averaging: Each Dictionary had 41 datasets. We averaged the values associated with same time value of datasets having same value of total momentum. Then using time reversal symmetry we averaged the 2pt values at time values (i) and (64-i) for i from 1 to 31. The values for time values 0 and 32 remain unchanged. This step is necessary to obtain a single plateau in the effective energy vs. time graph.
3.
Jackknife Analysis: We performed Jackknife resampling on the averaged data to calculate the Jackknife mean and standard deviation for each time step. This helps in estimating the uncertainty of the results.
4.
Energy Calculation and Identification of Plateau: We computed the ratio of two-point functions at consecutive time steps and took the natural logarithm to form the energy arrays. After plotting the vs graph we observed a plateau-like region, selected the range in terms of time step and mark it as plateau region.
5.
Calculation of Plateau Energy and Jackknife Error: Then we calculated the final value of effective energy () at that particular value of momentum using the following formula:
@@ 3 PROCEDURE
6.
Comparison of Dispersion Relation: Next we plot the dispersion relation using the results from various momentum analyses, applying Einstein’s special theory of relativity. In natural units: where is the lattice unit.
@@ 4 RESULTS and CONCLUSION
Following the already mentioned procedure, two representative plateaus are shown below for the momentum and (where the x and y component is 0) and momentum (where the resultant of x and y component is ).
plateau = 0.6430214 ± 0.0000028
plateau = 0.5456180 ± 0.0000017
Plateau fitted values for different combinations of momenta boost are given below in the table: Now we compare our computationally generated data (through Lattice QCD) with its theoretical counterpart using the dispersion relation
Momentum Configuration
Effective Energy
Error
0.2503897571908317
0.31812734594752745
2
0.3732724854969631
3
0.4247528925508162
0.463223332509825
5
0.5085321794457005
6
0.5438130645474166
8
0.5980768744904814
0.6356445374137877
10
0.6668746069233606
11
0.69971653807267
12
0.7251039340525522
13
0.747680677161704
14
0.7759013508398236
0.8289985937907043
17
0.854196899649006
18
0.8759581328257354
19
0.8991794712094623
20
0.91539787219901
21
0.9344281530391869
24
0.9972232821194545
1.00780980951629
27
1.0456348117309648
From the above plot we clearly see that our data clearly agrees with its theoretical counterpart so we can say that Lattice QCD can be one of the promising approach to probe the kaons in lower as well as for higher energy spectrum.
We thank Dr. Martha Constantinou for mentoring us throughout the project and the REYES program for allowing us to work on this project. We also thank Josh Miller and Isaac Anderson for their advice on technical aspects. M.C. acknowledges financial support from the U.S. Department of Energy, Office of Nuclear Physics, Early Career Award under Grant No. DE-SC0020405, and Grant No. DE-SC0025218. Interactions within the Quark-Gluon Tomography (QGT) Topical Collaboration, funded by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, with Award DE-SC0023646, have benefited this work. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award NP-ERCAP0027642, as well as NERSC award ALCC-ERCAP0030652. Also, some of the data used in this work have been produced using a computational award from Oak Ridge Leadership Computing Facility through the Summit Plus program (project: NPH160).